# Control of photoionization by resonant phase-locked pulse pairs

### v1

@@ I Introduction
What happens to atoms in strong fields? At low frequencies, where more than one photon is required for photoionization, there are two separate theoretical approximations: the strong-field approximation (SFA) [1] and the rotating-wave approximation (RWA) [2]. These approaches are complementary as they apply to non-resonant and resonant processes, respectively. In more extreme cases, atoms can be found in high-frequency superintense laser pulses, where one photon is enough for photoionization, but the forces from the laser field and the atomic potential are comparable [3]. This implies that neither SFA nor RWA are applicable and that another route is required. In this case, quasiclassical approaches have been proposed, based on the Wentzel–Kramers–Brillouin (WKB) approximation [4, 5], and by performing perturbative Coulomb corrections to SFA [6], but the simultaneous action of laser and atom can only be accounted for quantitatively by numerical propagation of the time-dependent Schrödinger equation (TDSE) [7, 8] or Dirac equation (TDDE) [9]. Interestingly, it has been found that all three strong-field regimes share the counterintuitive feature that ionization can be prevented (or reduced), a phenomenon commonly referred to as stabilization. Here, our focus will be on stabilization of atoms in strong fields, and, as proposed by Gavrila [3], we will distinguish between quasistationary stabilization (QS), which concerns decay rates obtained by Floquet-like theories for atoms in monochromatic fields, and dynamical stabilization (DS), which concerns time-dependent pulse shapes and atomic populations from TDSE (or TDDE) computations.
In this paper, we show how QS in two-photon resonant ionization (1+1), a phenomenon first predicted by Beers and Armstrong in 1975 [10], can be turned into a DS phenomenon using a sequence of two pulses, as schematically shown in Fig. 1.
@@
Our general ideas are inspired by prior works on alkali atoms by Wollenhaupt et al. from 2003 [11], where two optical laser pulses were used to create complex photoelectron patterns by coherent control of two dressed states. We study instead helium atoms in the extreme ultraviolet (XUV) regime, where the conditions required for QS are feasible [12], thanks to the development of seeded short-wavelength free-electron lasers (FELs) [13]. Recently, nonlinear interference phenomena [14] and quantum entanglement [15] have been detected for XUV-FEL pulses, and coherent control of ionization probabilities has been reported for chirped XUV-FEL pulses [16], opening the field of short-wavelength strong-coupling physics on ultrashort timescales. The latter phenomenon arises due to a DS process predicted by Saalmann et al. in 2018 [17], which can be interpreted using Beers and Armstrong’s QS regime, because a stabilized dressed state can be selected using chirped pulses by rapid adiabatic passage. We predict that coherent control of atomic stabilization can alternatively be achieved by using a sequence of two unchirped, phase-locked FEL pulses, provided that their intensities, pulse durations, polarizations, and photon energies are appropriately chosen, in accordance with the QS results reported in Ref. [12].
Our article is organized as follows: an overview of atomic stabilization phenomena is given in Sec. II, followed by an outline of our theory, with ionization rates extracted from quasienergy calculations in Sec. III, our results for time-dependent population dynamics and coherent-control of stabilization against photoionization in Sec. IV, and our conclusions in Sec. V. Atomic units are used unless otherwise stated: $\hbar=e=4\pi\epsilon_{0}=m_{e}=1$.
@@ II Overview of stabilization phenomena
In the following section, we present a summary of different stabilization phenomena that have been studied during the last decades. While the more general notion of bound states in the continuum goes back to early work on quantum mechanics by Wigner and von Neumann from 1929 [18], the subject has expanded to atoms in magnetic fields by Friedrich and Wintgen in 1984 [19], to optics with resonators and photonic structures, as reviewed by Koshelev et al. [20], and to scattering theory, as reviewed by Domcke [21]. The fact that atoms can stabilize in strong fields under vastly different conditions has also attracted much independent interest in the strong-field community. It is not possible here to account for all this development, but we aim to stress, by giving a few examples, that strong-field stabilization of atoms is a broad subject, beyond the superintense regime discussed by Gavrila [3]. After our brief historical account, we then proceed with the main results and discussions about the stabilization regime of interest in this work.
@@ II.2 High intensity and resonant frequency
The second theoretical line of strong-field ionization was initiated in 1975 by Beers, Armstrong, and Feneuille to describe resonant photoionization in the strong-coupling (SC) regime [10, 32]. In the SC regime, light-matter interactions can be understood within the RWA using dressed states [33, 2]. Beers and Armstrong [10] and Holt et al. [34] applied the RWA, using the resolvent operator method and effective Hamiltonian theory, to find that the intensity dependence of the ionization probability can be characterized in terms of three system parameters: the ionization rates of the essential atomic state, $\left|a\right\rangle$ and $\left|b\right\rangle$, denoted $\Gamma_{a}$ and $\Gamma_{b}$, respectively, and their mutual coupling coefficient, $q$. We note that $q$ accounts for both the direct transition between the essential states and higher-order virtual transitions via non-essential states (including the continuum). It was shown that atoms can photoionize in different ways, e.g., the atom may ionize primarily from the ground or excited atomic state. It was also proposed that quantum interference between these ionization pathways can play a role in the dynamics. In particular, it was predicted that there exist special parameters for which the ionization probability of the atom is bounded by some value $C$ such that $P_{\textrm{ion}}\leq C<1$ when the atom is subjected to a monochromatic field. For the resonant case one requires that $\Gamma_{a}\approx\Gamma_{b}$, which means that the atom ionizes with similar rates from both atomic states. One limitation of these early works was, however, that they only considered ionization to a single continuum [10, 34]. In 2023, the theory of Beers and Armstrong was generalized to multiple continua by Olofsson and Dahlström and applied to the experimentally relevant helium atom at the XUV transition: 1s2-1s2p [12]. It was found that dressed-atom stabilization was reached at an intensity of $\sim 10^{14}\,$W/cm2. In hindsight, the fact that atoms can photoionize at the same rate from both essential states in strong coupling was identified experimentally already in 2022 by Nandi et al. for the $1s^{2}-1s4p$ transition at $2\times 10^{13}$ W/cm2 using seeded XUV-FEL pulses with linear polarization [14]. More recently, experiments have been performed on the $1s^{2}-1s2p$ transition using circular polarized XUV-FEL pulses at intensities in the range of $10^{14}$ W/cm2 by Richter et al. [16]. Related stabilization effects, due to destructive interference of two leaky sources, have previously been found for atoms in strong magnetic fields [19] and for light propagation in photonic structures [20]. Dressed-atom stabilization physics corresponds to Keldysh parameters in the multiphoton regime, $\gamma_{\text{Keldysh}}\approx 10>1$.
@@ II.4 Stabilization of autoionizing resonances
Finally, we mention that autoionizing states in atoms, subjected to resonant strong laser fields, can lead to stabilization of one of the polaritonic branches (i.e., dressed states), as first predicted by Lambropoulos in 1981 [36]. Recently, such effects have been studied in experiments and theory developed further [37, 38, 39, 40]. The underlying physics that leads to stabilization of autoionizing polaritons is closely related to dressed-atom stabilization, since both phenomena arise due to interference of different ionization processes within the RWA. The physical origin of autoionization can be a static configuration interaction [36], induced dynamically by the same field that strongly couples the atoms [10], or by an additional laser field that couples to a continuum [41]. If the laser envelope shape can be coherently manipulated, the latter case opens up for independent and time-dependent control of autoionization processes and DS phenomena.
@@ III Theory
In the following subsections, we provide a brief account of the theory applied in this work. As is well known, the dynamics of two-level systems can be interpreted in terms of a torque vector, $\vec{\Omega}$, that rotates the pseudospin, $\vec{s}$, on the Bloch sphere according to [42, 43]
@@
where the torque vector can be written as $\vec{\Omega}=(\Omega\cos\phi,\Omega\sin\phi,\Delta)$ in Cartesian coordinates. More precisely, the Rabi frequency is calculated in the electric dipole form, $\Omega(t)=|\vec{d}\cdot\vec{E}_{0}(t)|$, the phase of the electric field is defined by $\vec{E}(t)=\vec{E}_{0}(t)\cos(\omega t-\phi)$, and the detuning of the field is taken relative to the atomic transition as $\Delta=\omega-\omega_{0}$. If we assume that $\Delta=0$, then the torque vector resides in the $xy$-plane and the pseudospin revolves on a great circle from the ground state to the excited state. The amount of revolution is determined by the area theorem, given by the pulse area, $\theta=\int dt\Omega(t)$.
In this work, we will consider quantum coherent control with two sequential resonant pulses (a pump pulse and a probe pulse) that have variable relative field phase, $\phi$. Assuming that the pump pulse is a $\theta_{\mathrm{pump}}=\pi/2$ pulse with $\phi_{\mathrm{pump}}=0$, the field phase of the probe pulse can be chosen to make its corresponding torque vector parallel with the pumped pseudospin by $\phi_{\mathrm{probe}}=\pm\pi/2$. The probe pulse will then not drive any further dynamics on the Bloch sphere, as follows from the cross product in Eq. (1). There are two such stationary cases, corresponding to parallel and anti-parallel orientation of torque and pseudospin. In other words, the pumped state
$\left|\Psi_{\mathrm{pump}}\right\rangle=\frac{1}{\sqrt{2}}(\left|a\right\rangle-i\left|b\right\rangle),$ (3)
can be projected onto one of the two phase-locked dressed states of the probe pulse,
$\left|\pm_{\phi}\right\rangle=\frac{1}{\sqrt{2}}\left(\left|a\right\rangle\pm e^{i\phi}\left|b\right\rangle\right),$ (4)
where $\phi$ is the field phase of the probe pulse. The population of the two possible eigenstates, $\left|\pm_{\phi}\right\rangle$ with eigenvalues $\pm\Omega/2$, is controlled by the sign of the probe field phase, $\phi=\pm\pi/2$.
III.1 Extracting ionization rates from quasienergies
@@
where $H_{0}$ is the field-free atomic Hamiltonian, $D=\mathrm{diag}(-n,\dots,0,\dots,m)$, $z$ is the $z$-component of the atomic dipole operator, and $F$ is a matrix with ones on the sub- and superdiagonals. For circular polarization in the $xy$-plane, the co-rotating frame formulation leads to the following Hamiltonian
@@
III.2 Stabilization of dressed states
@@
with $\beta=27/8$ a.u.-1 [64]. While the SAE approach qualitatively reproduces the minimum in ionization rate from the $\left|+\right\rangle$ state, its position in intensity is higher by a factor of 6, compared to the two-electron calculation (CIS). Surprisingly, this shows that many-body effects are essential for a quantitative description of the dressed-atom stabilization phenomenon. We attribute a part of the difference between the CIS and SAE results to the fact that the ground state orbital is doubly occupied with spin singlet symmetry. In TDCIS, this leads to an extra factor of $\sqrt{2}$ in the equations of motion for the dipole excitation from the ground state to the excited states [65], increasing the corresponding ionization rate by a factor of $2$. The same phenomenon also leads to $\sqrt{2}$ times faster Rabi oscillations in TDCIS compared to SAE, but it does not alter the ionization rate from the excited state. We refer the reader to the Appendix for a more detailed numerical study of the gauge invariance of the ionization rates in the full calculation and its dependence on state truncation.
III.3 Time-dependent approaches

### v2

# Coherent control of ionization via stabilization by resonant pulse pairs

@@ I Introduction
What happens to atoms in strong fields? At low frequencies, where more than one photon is required for photoionization, there are two separate theoretical approximations: the strong-field approximation (SFA) [4] and the rotating-wave approximation (RWA) [14]. These approaches are complementary as they apply to non-resonant and resonant processes, respectively. In more extreme cases, atoms can be found in high-frequency superintense laser pulses, where one photon is enough for photoionization, but the forces from the laser field and the atomic potential are comparable [29]. This implies that neither SFA nor RWA are applicable and that another route is required. In this case, quasiclassical approaches have been proposed, based on the Wentzel–Kramers–Brillouin (WKB) approximation [23, 39], and by performing perturbative Coulomb corrections to SFA [58], but the simultaneous action of laser and atom can only be accounted for quantitatively by numerical propagation of the time-dependent Schrödinger equation (TDSE) [20, 80] or Dirac equation (TDDE) [42]. Interestingly, it has been found that all three strong-field regimes share the counterintuitive feature that ionization can be prevented (or reduced), a phenomenon commonly referred to as stabilization. Here, we focus on stabilization of atoms in strong fields, and, as proposed by Gavrila [29], we will distinguish between quasistationary stabilization (QS), which concerns decay rates obtained by Floquet-like theories for atoms in monochromatic fields, and dynamical stabilization (DS), which concerns time-dependent pulse shapes and atomic populations from TDSE (or TDDE) computations. In the QS sense, stabilization is characterized by a decrease or oscillatory behavior of ionization rates with increasing intensity [29]. On the other hand, DS is characterized by the behavior of the ionization probability at the end of a pulse sequence, if the ionization probability levels out or decreases with an increase in the intensity [29].
In this paper, we show how QS in two-photon resonant ionization (1+1), a phenomenon first predicted by Beers and Armstrong in 1975 [6], can be used to interpret a DS phenomenon using a sequence of two pulses, as schematically shown in Fig. 1.
@@
Our general ideas are inspired by prior works on alkali atoms by Wollenhaupt et al. from 2003 [74], where two optical laser pulses were used to create complex photoelectron patterns by applying coherent control with two dressed states. We study, instead, helium atoms in the extreme ultraviolet (XUV) regime, where the conditions required for QS are feasible [55], thanks to the development of seeded short-wavelength free-electron lasers (FELs) [2]. Recently, nonlinear interference phenomena [51] and quantum entanglement [52] have been detected for XUV-FEL pulses, and coherent control of ionization probabilities has been reported for chirped XUV-FEL pulses [60], opening the field of short-wavelength strong-coupling physics on ultrashort timescales. The latter phenomenon arises due to a DS process predicted by Saalmann et al. in 2018 [61], which can be interpreted using Beers and Armstrong’s QS regime, because a stabilized dressed state can be selected using chirped pulses by rapid adiabatic passage. We predict that coherent control of atomic stabilization can alternatively be achieved by using a sequence of two unchirped, phase-locked FEL pulses, provided that their intensities, pulse durations, polarizations, and photon energies are appropriately chosen, in accordance with the QS results reported in Ref. [55].
Our article is organized as follows: an overview of atomic stabilization phenomena is given in Sec. II, followed by an outline of our theory, with ionization rates extracted from quasienergy calculations in Sec. III, our results for time-dependent population dynamics and coherent-control of stabilization against photoionization in Sec. IV, and our conclusions in Sec. V. There are also two appendices that discuss gauge invariance for restricted basis set calculations, App. A, and how dipole moments differ between configuration interaction singles and single active electron models, App. B. Atomic units are used unless otherwise stated: $\hbar=e=4\pi\epsilon_{0}=m_{e}=1$.
@@ II Overview of stabilization phenomena
In the following section, we present a summary of different stabilization phenomena, both QS and DS, that have been studied during the last decades. While the more general notion of bound states in the continuum goes back to early work on quantum mechanics by Wigner and von Neumann from 1929 [73], the subject has expanded to atoms in magnetic fields by Friedrich and Wintgen in 1984 [27], to optics with resonators and photonic structures, as reviewed by Koshelev et al. [43], and to scattering theory, as reviewed by Domcke [17]. The fact that atoms can stabilize in strong fields under vastly different conditions has also attracted much independent interest in the strong-field community. It is not possible here to account for all this development, but we aim to stress, by giving a few examples, that strong-field stabilization of atoms is a broad subject, beyond the superintense regime discussed by Gavrila [29]. After our brief historical account, we then proceed with the main results and discussions about the stabilization regime of interest in this work.
@@ II.2 High intensity and resonant frequency
The second theoretical line of strong-field ionization was initiated in 1975 by Beers, Armstrong, and Feneuille to describe resonant photoionization in the strong-coupling (SC) regime [6, 5]. In the SC regime, light-matter interactions can be understood within the RWA using dressed states [13, 14]. Beers and Armstrong [6] and Holt et al. [35] used effective Hamiltonian theory to find that the intensity dependence of the ionization probability can be characterized in terms of three system parameters: the ionization rates of the essential atomic state, $\left|a\right\rangle$ and $\left|b\right\rangle$, denoted $\Gamma_{a}$ and $\Gamma_{b}$, respectively, and their mutual coupling coefficient, $q$. We note that $q$ accounts for both the direct transition between the essential states and higher-order virtual transitions via non-essential states (including the continuum). It was shown that atoms can photoionize in different ways, e.g., the atom may ionize primarily from the ground or excited atomic state. It was also proposed that quantum interference between these ionization pathways can play a role in the dynamics. In particular, it was predicted that there exist special parameters for which the ionization probability of the atom is bounded by some value $C$ such that $P_{\textrm{ion}}\leq C<1$ when the atom is subjected to a monochromatic field. For the resonant case one requires that $\Gamma_{a}\approx\Gamma_{b}$, which means that the atom ionizes with similar rates from both atomic states. One limitation of these early works was, however, that they only considered ionization to a single continuum [6, 35]. In 2023, the theory of Beers and Armstrong was generalized to multiple continua by Olofsson and Dahlström and applied to the experimentally relevant helium atom at the XUV transition: 1s2-1s2p [55]. It was found that dressed-atom stabilization was reached at an intensity of $\sim 10^{14}\,$W/cm2. In hindsight, the fact that atoms can photoionize at the same rate from both essential states in strong coupling was identified experimentally already in 2022 by Nandi et al. for the $1s^{2}-1s4p$ transition at $2\times 10^{13}$ W/cm2 using seeded XUV-FEL pulses with linear polarization [51]. More recently, experiments have been performed on the $1s^{2}-1s2p$ transition using circular polarized XUV-FEL pulses at intensities in the range of $10^{14}$ W/cm2 by Richter et al. [60]. Related stabilization effects, due to destructive interference of two leaky sources, have previously been found for atoms in strong magnetic fields [27] and for light propagation in photonic structures [43]. Dressed-atom stabilization physics corresponds to Keldysh parameters in the multiphoton regime, $\gamma_{\text{Keldysh}}\approx 10>1$.
@@ II.4 Stabilization of autoionizing resonances
Finally, we mention that autoionizing states in atoms, subjected to resonant strong laser fields, can lead to stabilization of one of the polaritonic branches (i.e., dressed states), as first predicted by Lambropoulos in 1981 [46]. Recently, such effects have been studied in experiments and theory developed further [44, 33, 76, 9]. The underlying physics that leads to stabilization of autoionizing polaritons is closely related to dressed-atom stabilization, since both phenomena arise due to the interference of different ionization processes within the RWA. The physical origin of autoionization can be a static configuration interaction [46], induced dynamically by the same field that strongly couples the atoms [6], or by an additional laser field that couples to a continuum [15]. If the laser envelope shape can be coherently manipulated, the latter case opens up for independent and time-dependent control of autoionization processes and DS phenomena.
@@ III Theory
In the following subsections, we provide a brief account of the theory applied in this work.
III.1 Form of the laser pulses
In this work, the total laser vector potential is given by
$A(t)=A_{\textrm{pump}}(t)+A_{\textrm{probe}}(t),$ (1)
and the electric field is $E(t)=-\dot{A}(t)$. The pump and probe pulses are defined by
$A_{\textrm{i}}(t)=A_{0}f(t)[A_{x}(t),A_{y}(t),A_{z}(t)],$ (2)
where $A_{0}$ is the amplitude of the vector potential and $f(t)$ is the pulse envelope
$f(t)=\begin{cases}\cos^{2}\left[\frac{\pi(t-t_{0})}{\tau}\right],&|t-t_{0}|<\frac{\tau}{2}\\ 0,&|t-t_{0}|\geq\frac{\tau}{2}.\end{cases}$ (3)
In Eq. (3), $t_{0}$ and $\tau$ correspond to the pulse delay and pulse duration (foot-to-foot), respectively. In the case of linear polarization, the field components are chosen as
$\begin{split}&A_{x}(t)=0,\\ &A_{y}(t)=0,\\ &A_{z}(t)=\sin\left[\omega(t-t_{0})-\varphi\right],\end{split}$ (4)
where $\varphi$ is the carrier–envelope phase (CEP). For circular polarization the components are given by
$\begin{split}&A_{x}(t)=\frac{1}{\sqrt{2}}\sin\left[\omega(t-t_{0})-\varphi\right],\\ &A_{y}(t)=-\frac{1}{\sqrt{2}}\cos\left[\omega(t-t_{0})-\varphi\right],\\ &A_{z}(t)=0.\end{split}$ (5)
The probe pulse is delayed by a fixed amount $t_{0}>\tau$ with a variable CEP $\varphi\in[0,2\pi]$, while the pump pulse is fixed to give a zero reference at $t_{0}^{\mathrm{pump}}=0$ with $\varphi^{\mathrm{pump}}=0$. In this configuration, we define the field phase of the probe field
$\phi=\omega t_{0}+\varphi.$ (6)
III.2 Two-level system dynamics
We will now provide a brief discussion of the dynamics of a two-level system that is relevant for our proposed scheme. For simplicity, we will work with linear polarization, but the ideas also apply to circular polarization. In this case, the electric field of a single pulse takes the form $E(t)=\-\vec{E}_{0}(t)\cos(\omega t-\phi)$, where $E_{0}(t)=\omega A_{0}f(t)$ and we have assumed that the envelope varies slowly enough for its derivative to be negligible. The atomic states that make up the two-level system are the ground, $\left|a\right\rangle$, and the excited state, $\left|b\right\rangle$.
As is well known, the dynamics of two-level systems can be interpreted in terms of a torque vector, $\vec{\Omega}$, that rotates the pseudospin, $\vec{s}$, on the Bloch sphere according to [25, 3]
@@ III.2 Two-level system dynamics
where the torque vector can be written as $\vec{\Omega}=({\color[rgb]{0,0,0}-}\Omega\cos\phi,{\color[rgb]{0,0,0}-}\Omega\sin\phi,\Delta)$ in Cartesian coordinates and our choice of field parametrization. More precisely, the Rabi frequency is calculated in the electric dipole form, $\Omega(t)=|z_{ba}E_{0}(t)|$, where $z_{ba}=\left\langle b\right|z\left|z\right\rangle$ is the dipole matrix element between the ground and excited states, and the detuning of the field is taken relative to the atomic transition as $\Delta=\omega-\omega_{0}$. If we assume that $\Delta=0$, then the torque vector resides in the $xy$-plane and the pseudospin revolves on a great circle from the ground state to the excited state. The amount of revolution is determined by the area theorem, given by the pulse area, $\theta=\int dt\Omega(t)$. The eigenvectors of the Hamiltonian on the right-hand side of Eq. (8) have the form
$\left|\pm\right\rangle=\frac{1}{\sqrt{2}}\left(\left|a\right\rangle{\color[rgb]{0,0,0}\mp}e^{i\phi}\left|b\right\rangle\right),$ (9)
and are known as dressed states. Their eigenvalues are given by $\pm\Omega/2$.
Assuming that the pump pulse is a $\theta_{\mathrm{pump}}=\pi/2$ pulse with $\phi_{\mathrm{pump}}=0$, the pseudospin after the pump pulse will lie in the $xy$-plane. The field phase of the probe pulse can then be chosen to make its corresponding torque vector parallel with the pumped pseudospin by $\phi_{\mathrm{probe}}=\pm\pi/2$. The probe pulse will then not drive any further dynamics on the Bloch sphere, as follows from the cross product in Eq. (7). There are two such stationary cases, corresponding to parallel and anti-parallel orientation of torque and pseudospin. This corresponds to the situation where the state after the pump pulse, $\left|\Psi_{\mathrm{pump}}\right\rangle$, is one of the dressed states of the probe pulse. In general, $\left|\Psi_{\mathrm{pump}}\right\rangle$ will be a superposition of $\left|\pm\right\rangle$, and the relative population can be controlled by changing $\phi_{\mathrm{probe}}$.
III.3 Extracting ionization rates from quasienergies
@@
where $H_{0}$ is the field-free atomic Hamiltonian, $\mathbb{1}$ is the identiy matrix, $D=\mathrm{diag}(-n,\dots,0,\dots,m)$, $z$ is the $z$-component of the atomic dipole operator, $F$ is a matrix with ones on the sub- and superdiagonals, and $\otimes$ denotes the tensor product of the atomic Hilbert space with the Floquet basis. For circular polarization in the $xy$-plane, the co-rotating frame formulation leads to the following Hamiltonian
@@
III.4 Stabilization of dressed states
@@
with $\beta=27/8$ a.u.-1 [34]. While the SAE approach qualitatively reproduces the minimum in the ionization rate from the $\left|+\right\rangle$ state, its position in intensity is higher by a factor of 6, compared to the two-electron calculation (CIS). Surprisingly, this shows that many-body effects are essential for a quantitative description of the dressed-atom stabilization phenomenon. We attribute a part of the difference between the CIS and SAE results to the fact that the ground state orbital is doubly occupied with spin singlet symmetry. In TDCIS, this leads to an extra factor of $\sqrt{2}$ in the equations of motion for the dipole excitation from the ground state to the excited states [32], see App. B for a derivation of this fact, increasing the corresponding ionization rate by a factor of $2$. This is a simple statistical factor due to the fact that either of the two electrons (spin up or down) can be ionized and contribute to the total rate. The same phenomenon also leads to $\sqrt{2}$ times faster Rabi oscillations in TDCIS compared to SAE, but it does not alter the ionization rate from the excited state. Another potential source of discrepancy between the models could be the shape of the radial orbitals, which will affect the values of the dipole matrix elements. We refer the reader to App. A for a more detailed numerical study of the gauge invariance of the ionization rates in the full calculation and its dependence on state truncation.
III.5 Time-dependent approaches